Problem 3-2 National Scan, Inc., sells radlo frequency Inventory tags. Monthly sales for a seven-month period were as follows: Sales Month(ee)Units Feb Mar Apr May Jun Jul Aug 2e 23 25 18 28 30 b. Forecast September sales volume using each of the folloWng: (1 A linear trend equation.(Round your Intermedlate calculatlons and final answer to 2 declmal places.) Yt thousands (2) A five-month moVing average. (Round your answer to 2 declmal places.) Moving averagehousands (3) Exponentlal smoothing with a smoothing constant equal to .20, assuming a March forecast of 18(000). (Round your Intermediete forecast values and final answer to 2 declmel places Forecast thousands (4) The nalve approach. Nalve approach thousands (5) A weighted average using .60 for August, 20 for July, and .20 for June. (Round your answer to 2 declmal places.) Weighted average thousands
Answer #1

The data given is as under:

Month Sales (000) Units # 1 Feb 2 Mar 20 23 3 Apr 4 May Jun 6 July Aug 18 28 30

A) LINEAR TREND EQUATION

A linear trend equation can be computed using MS excel, plotting a graph and adding a trendline to it.

National Scan Inc. 35 30 25 20 15 10 30 y 1.8214 x + 14.286 25 23 20 18 Apr May Jun July Aug Feb Mar

The linear trend equation is, Y = 1.8214 X + 14.286

For the month of September (x=8), the forecast would be

= (1.8214 x 8) + 14.286 = 28.8572 Units = 29 Units approx.

B) FIVE MONTH MOVING AVERAGE

The average of last five months (Apr-Aug) will be taken as the new forecast for September

= (7+25+18+28+30)/5 = 21.6 units

C) EXPONENTIAL SMOOTHING

Smoothing constant = 0.20

March Forecast = 18 units

The formula for exponential smoothing forecast is-

Forecast (next month) = Forecast (current month) + [Smoothing constant x )Actual(current month) - Forecast(current month) ) ]

The forecasted values are as under,

Actual Sales Forecasted Month Feb Mar Apr May Jun July Aug Sep (000) Unitssales 20 18.000 19.000 16.600 18.280 18.224 20.179

The forecast for the month of september is = 22.143 units

D) NAIVE APPROACH gives the last month actual sales as the new forecast.

The forecast for the month of September = Actual sales in Aug = 30 Units

E) WEIGHTED AVERAGE APPROACH

Aug = 0.60, July = 0.20, June = 0.20

Forecast for Sep = (Wt. of Aug X Actual sales in Aug) + (Wt. of July X Actual sales in July) + (Wt. of June X Actual sales in June)

= (0.60 x 30) + (0.20 x 28) + (0.20 x 18) = 27.2 Units